Calculate the temperature of a town situated on the top of a mountain at the height 600m above sea level, given that the sea level temperature is 37 degrees
1
Expert's answer
2020-10-15T06:37:22-0400

Every 1000 m (1 km), the temperature drops for 6°C. Therefore, 600 meters correspond to


"t=\frac{600}{1000}\cdot6=3.6°\text{C}."

The temperature of air near at the height of the town is


"t_t=37-3.6=33.4°\text{C}."